OpenStudy (anonymous):

stabdard form of a line passing through (-2,7) and a slope of -3/5

OpenStudy (cj49):

y - y1 = m(x -x1)

OpenStudy (anonymous):

You can use the formula $$ (y - y_1) = m(x-x_1) $$ Plug in -2 for x1, 7 for y1, and (-3/5) for m. Carry over y1 to the other side, and simplify.

OpenStudy (anonymous):

thank yall! :)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 7}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\cfrac{3}{5} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and move everything to the left}\\ \qquad \uparrow\\ \textit{point-slope form}\)